Ron's age is two-thirds of Mark's age. Mark's age is three-fifths of Sara's age. Sara is 9 years older than Ron. How old is Mark?
Answer by mananth(16946)

Sara x years
Mark (3/5)x
Ron (2/3) * (3/5)x=(2x/5)
x= (2x/5)+9
x-2x/5 = 9
3x/5=9
3x=45
x= 15 Ron's age
Mark's age (3/5)x = (3/5)*15= 9 years